Problem #61 on p. 150 Chapter 2.1 reads - 2/3 - (- 5/8). We've done these before so I don't really know what I keep doing wrong that I keep getting so stuck on these. I know that you change the sign and add the opposite so it would then look like this - 2/3 + 5/8, and I know you must choose a common denominator as well which in this case would be 24. I also know the answer must be a negative because - 2/3 is the greater fraction...meaning if you have 5/8's and you owe 2/3's you are still going to owe something. What I keep wanting to do is -2 + 5 which gives you 3 but that is a positive number so I know it can't be right. I'm getting really frustrated at the fact that I suddenly can't seem to once again recall how to subtract fractions properly. I'm really stuck at this point and could use another reminder and perhaps an easier way to recall how exactly to do this so that I will hopefully be able to keep it in mind and just do them right without panicking or getting stuck from here on out. Can someone help? Thanks a lot!
Answer by tutorcecilia(2144)

To add or subtract fractions, remember this rule and follow the pattern:
a/b + c/d = Cross-multiply and than multiply the denominators:
.
(a)(d) + (b)(c)/(b)(d)
.
The same rule applies for subtraction:
a/b - c/d = (a)(d) - (b)(c)/(b)(d)
.
You are correct regarding adding integers (negative and positive numbers):
To add integers with the same sign, add the absolute values and take the common sign. Example: 2+5= 7 or -2+(-5)=-7
.
To add integers with different signs, subtract the absolute values and take the sign of the larger number. for example: -2+5 = 3 or 2-5 = -3
.
To subtract numbers with either the same or different signs, change the subtraction to addition and change the sign of the second number. Than follow whichever rules of addition apply.
Solving - 2/3 + 5/8
-2/3 + 5/8
Cross-multiply and multiply the denominators:
(-2)(8) + (3)(5)/(3)(8) = -16+15/24 = -1/24
I am also available for one-on-one e-mail tutoring. Just send problems to tutorcecilia@yahoo.com. Thanks and good luck!.